In 10 years,A will be twice as old as B was 10 Years ago.If A is now 9 years older than B,the present age of B is?
Answer by ankor@dixie-net.com(22740)

In 10 years, A will be twice as old as B was 10 Years ago.
a + 10 = 2(b-10)
a + 10 = 2b - 20
a = 2b - 20 -10
a = 2b - 30
:
If A is now 9 years older than B,
a = b + 9
:
the present age of B is?
replace a with (2b-30)
2b - 30 = b + 9
2b - b = 9 + 30
b = 39 yrs old
:
:
:
Check this in the statement: (a: 39+9 = 48)
"In 10 years, A will be twice as old as B was 10 Years ago."
48 + 10 = 2(39-10)
58 = 2(29)